OpenStudy (anonymous):

what is the solution of the problem x-2/4=x+10/10

OpenStudy (anonymous):

\[\frac{ x-2 }{ 4 }= \frac{ x+10 }{ 10 }\]

OpenStudy (anonymous):

this is how the problem looks

OpenStudy (anonymous):

multiply both sides by 10 and multiply both sides by 4

OpenStudy (acxbox22):

cross multiply and solve the equation

OpenStudy (acxbox22):

\[4(x+10)=10(x-2)\]

OpenStudy (anonymous):

4x+40=10x-20 6x=60 x=10

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

your welcome:)

OpenStudy (anonymous):

i need help on another question too i can't cross multiply

OpenStudy (anonymous):

\[\frac{ 8 }{ x+2 }= \frac{ 3 }{ x-1 }\]

OpenStudy (acxbox22):

csscgirl, refrain from giving away answers but guiding them through it.

OpenStudy (anonymous):

wait can i cross multiply in this one?

OpenStudy (acxbox22):

yes

OpenStudy (anonymous):

so it will be 8(x-1)=3(x+2)

OpenStudy (acxbox22):

correct

OpenStudy (anonymous):

i'm getting 5x=7 so far i'm i correct?

OpenStudy (anonymous):

which will result in x=1.4

OpenStudy (acxbox22):

no, you will get 5x=14

OpenStudy (anonymous):

but you divide 7/5 which results in 1.4 which is one of the options for my hw

OpenStudy (acxbox22):

from my answer i got 2.8 and it works in your equation

OpenStudy (anonymous):

oh how did you get 14?

OpenStudy (acxbox22):

8(x-1)=3(x+2) 8x-8=3x+6 8x=3x+14 5x=14

OpenStudy (anonymous):

ok i see what i did wrong, i didnt multiply 8 by 1 i just left it as 1.

OpenStudy (acxbox22):

cool

OpenStudy (anonymous):

thank you @acxbox22

OpenStudy (acxbox22):

np